OpenStudy (anonymous):

Evaluate w/o using a calulator: sec (11 pi/6) My answer was (2 sqrt3/3) Can someone check if it's right?

OpenStudy (anonymous):

sec (11π / 6) is the same as 1 / cos(11π / 6)

OpenStudy (anonymous):

sec(11pi/6) = 1/cos(11pi/6) cos(11pi/6) we can see from the graph that cos(2pi - x) = cos(x) cos(11pi/6) = cos(2pi - pi/6) = cos(pi/6) = sqrt(3)/2

OpenStudy (anonymous):

so sec(11pi/6) = 1/(sqrt(3)/2) = 2/sqrt(3) = 2sqrt(3)/3

OpenStudy (anonymous):

yes you are correct

OpenStudy (anonymous):

\[1 / \cos(11π / 6) = 1 / \sqrt{3}/2\] \[= 2\sqrt{3}/3\]

OpenStudy (anonymous):

Thanks smart people lol i thought i had it wrong. :)